OpenStudy (anonymous):

having serious problems finding the answer to this.. let f(x)= 3x^4-4x^3-12x^2+14 a.) use the first derivative to find exact values of all critical points and identify each critical point as a local minima, local maxima, or neither b) use the second derivative to find exact values of all inflection points of f(x)

OpenStudy (anonymous):

\[f \prime(x)= 12x^{3}-12x ^{2}-24x =12x(x ^{2}-x-2) =12x(x+1)(x-2)=0\] The values of x are: 0.-1,2 \[f \prime \prime (x) = 36x ^{2}-24x-24=0 \to 3x ^{2}-2x-2=0 \]